what is the equation of y^2= 1/5x when translated up 3 units and left 2 units
Answer by lwsshak3(11628)

what is the equation of y^2= 1/5x when translated up 3 units and left 2 units
y^2= x/5
y=±√(x/5)
translated up 3 units and left 2 units
y=±√[(x+2)/5]+3